SA: Lee
SO: Veronica
Others: Fred, Antoine, Thor, Lizette
------------------------------------------------------------------------
Summary
------------------------------------------------------------------------
The thunderstorms from earlier in the day eventually cleared out and left us with just over half the night to observe with clear skies and good seeing. We completed 6 high priority dark moon blocks.
Science data taken for:
2024-2-LSP-001 P0 RSS (Followup of Transients)
2025-2-MLT-006 P1 RSS (Study of galaxy properties of galaxy group HIPASS 1059-09)
2025-2-MLT-009 P1 RSS (x2) (Spectroscopy of ultra-diffuse galaxies)
2021-2-MLT-004 P1 RSS (Spectropolarimetric Monitoring of eta Car)
2025-2-SCI-023 P1 RSS (Measuring the orbit and disc of a neutron star X-ray binary)
------------------------------------------------------------------------
No comments:
Post a Comment